OpenStudy (mendicant_bias):

(Sturm-Liouville) I'm having trouble understanding how in a given problem they get an eigenfunction that they did. I understand how they got their eigenvalue. More info below.

OpenStudy (mendicant_bias):

In page 13 of the following PDF, problem III, I get how they got the eigenvalue to be \[\lambda=\bigg(\frac{n \pi}{l}\bigg)^2\] But I do not understand why they got the sine function as their eigenfunction. Isn't the eigenfunction supposed to be the result of taking your eigenvalue and plugging it back into your original, general solution? This would yield \[y_n=\cos\bigg(\frac{n \pi x}{l}\bigg)\]To me.

OpenStudy (mendicant_bias):

I'm thinking this must just be a typo or something, but does what I'm saying make sense?

OpenStudy (mendicant_bias):

(?)

OpenStudy (mendicant_bias):

I'm literally just asking if this is a typo, could somebody who knows this better than me help me out? I'm 90% sure this is just a mistake, but virtually all it takes is a "yes" or "no" from someone who is familiar with this stuff.

OpenStudy (mendicant_bias):

@mathmath333

OpenStudy (mendicant_bias):

Thanks, someone just told me that I didn't link the PDF. Somebody mentioning that prior would've been great. http://texas.math.ttu.edu/~gilliam/ttu/ode_pde_pdf/Ch5.pdf

OpenStudy (dan815):

wait wahts the confusion its right

OpenStudy (dan815):

notice the eigen values are just different form from 2 and 3

OpenStudy (mendicant_bias):

But where is the additional pi/2 coming from? The argument of the original function being \[y_h=c_1\sin(\alpha x)+c_2\cos(\alpha x)\]Where c_1=0, evaluating the original function at the eigenfunction gives \[y_n =\cos\bigg(\frac{n \pi x}{l}\bigg)\]

OpenStudy (mendicant_bias):

*evaluating the original function at the eigenvalue

OpenStudy (dan815):

u cancelled out the wrong part

OpenStudy (dan815):

c2=0

OpenStudy (mendicant_bias):

\[y_h=c_1\sin(\alpha x)+c_2\cos(\alpha x)\]\[y'_h=c_1 \alpha \cos(\alpha x)-c_2 \alpha \sin(\alpha x)\] This is for Part III of the problem, not Parts I or II.

OpenStudy (mendicant_bias):

\[y'(0)-c_1 \alpha \cos(0)+c_2 \alpha \sin(0)=0, --> c_1=0.\]

OpenStudy (mendicant_bias):

^Equals sign, not minus, and minus, not plus.

OpenStudy (dan815):

they are saying b=0 so u have to start with y'(l) first

OpenStudy (dan815):

y'(0) doesnt give us any useful information in this case

OpenStudy (mendicant_bias):

I don't understand what the issues is, I'll put it like this: We agree that the eigenvalue is n^pi^2/l^2, right? We do, don't we?

OpenStudy (mendicant_bias):

The way we find an eigenfunction is by taking the eigenvalue and substituting it into the original function. If b (or c_1)=0, then for your original function, you have \[y+h=c_1\cos(\alpha x)\] Right, @dan815 ???

OpenStudy (mendicant_bias):

*y_h

OpenStudy (mendicant_bias):

\[y_n=c_1\cos(\alpha x)=c_1 \cos\bigg(\frac{n \pi x}{l}\bigg)\]

OpenStudy (mendicant_bias):

@bibby

OpenStudy (mendicant_bias):

We agree. We agree, right? We agree that \[\lambda =\frac{n^2\pi^2}{l^2}\]

OpenStudy (dan815):

your method is not wrong, its just in this question b=0

OpenStudy (mendicant_bias):

The eigenfunction is determined by taking the eigenvalue-which we both agree is the above-and plugging it into the general solution, yes? Yes, we both agree that b is zero, I'm not disputing that, at all, whatsoever, not at all, dan, that's not what I'm confused about.

OpenStudy (dan815):

ok so

OpenStudy (mendicant_bias):

If b is zero, then the original solution must be \[y_h=c_1\cos(\alpha x)\] Right?

OpenStudy (dan815):

right

OpenStudy (dan815):

so now u do y'(l)

OpenStudy (dan815):

oh i see your ur problem u are confused about their final answer is in sin

OpenStudy (dan815):

thats obviously a mistake xD

OpenStudy (mendicant_bias):

Taking the agreed upon eigenvalue-but that doesn't matter, we already know about and agree about the eigenvalue, we get our eigenfunction by substituting our general solution, no derivatives or anything, right? If you have your eigenfunction-HWSD;KJSDFKL;SL;

OpenStudy (mendicant_bias):

Alright, I'm done, goodnight. Thank you, Dan.

OpenStudy (dan815):

lol np

OpenStudy (mendicant_bias):

Yeah, I just kept earlier over and over asking if it was a typo, that's what I was concerned about.

OpenStudy (dan815):

trying to do too many questions at once

OpenStudy (dan815):

im just skimming over comments

OpenStudy (mendicant_bias):

No worries. thanks again.